# Quantum Neural Networks under Depolarization Noise: Exploring White-Box Attacks and Defenses

### 29 Nov 2023 v1

@@ 1 Introduction
The field of quantum computing has shown fast advancements in recent years, providing the possibility to speed up speficic tasks. Hereby a variety of algorithms exists such as Shor’s algorithm to factor primes [30], Grover’s search to find an element in an unsorted list in $\mathcal{O}(\sqrt{n})$ [10] or the HHL-Algorithm to efficiently solve systems of linear equations [11]. In this context, quantum machine learning (QML) focuses onto providing the properties and speedups of quantum computing in the field of machine learning [33, 32, 28, 1]. First coined by Kak [13], the term quantum neural network (QNN) emerged as an ambitious effort to merge observations from the domain of neuroscience with the intriguing features inherent to quantum computations. Essentially, QNNs represent a specific type of Quantum Machine Learning (QML) models. Inspired by the layered approach of classical neural networks.
As their classical counterpart, Lu et al. [19] highlighted that QNNs are not immune to adversarial attacks. Adversarial attacks involve carefully designed disturbances that can mislead machine learning models, causing them to produce incorrect predictions. Such attacks have long been a significant issue in traditional machine learning [2, 31, 9, 20]. The importance of investigating adversarial attacks on QNNs is reinforced by the no free lunch theorem [24]. This principle suggests that there is not a single model or algorithm universally optimal for all tasks, hinting at inherent vulnerabilities also in quantum-based models. Hence, finding these weak spots is crucial for the resilience and protection of quantum machine learning systems. A further obstacle is that current quantum hardware used to execute these models suffers from noise and decoherence times, which is the reason why Preskill [25] coined the term NISQ-Era for the current state in quantum computing. This means, that due the noise, limited connectivity and decoherence times, inherent in quantum devices, current algorithms can only be executed on smaller toy problems, providing a proof of concept, but suffering from not yet being applicable to practical problems. In this work, we explore the effect of noise, specifically depolarization noise, on the robustness of Quantum Neural Networks (QNNs), aiming to minimize compromises on accuracy and training data. We assumed the structure in Figure 1 as the most realistic scenario, so the input is perturbed on the classical side and then encoded on the quantum circuit. Furthermore, our research builds on the theoretical insights of [6], which suggest that QNNs inherently possess an upper bound of robustness relative to the device’s depolarization noise. This upper bound arises because depolarizing noise, viewed as a stochastic Pauli-channel, effectively scrambles the state of the quantum device, thus blurring information. This blurring leads to a natural tradeoff between accuracy and robustness. We evaluate the effectiveness of this phenomenon by adding multiple depolarizing noise channels to a quantum neural network and comparing its performance with other adversarial training techniques commonly used in classical settings.
[…]
@@
Contrary to the findings of [6], our results indicate that depolarization noise does not improve adversarial robustness in multi-class classifiers. Additionally, we observed that increasing the number of classes diminishes both accuracy and overall robustness, with depolarization noise offering no significant enhancement.
@@ 2.1 Quantum information and Noise
$\psi$ $=$ $\alpha\ket{0}+\beta\ket{1}=\alpha\begin{pmatrix}1\\ 0\end{pmatrix}+\beta\begin{pmatrix}0\\ 1\end{pmatrix},$ (5)
with $~{}{|\alpha|}^{2}+{|\beta|}^{2}=1.$ (6)
Unlike classical computers that use bits with values of either 0 or 1, quantum computers can simultaneously represent both states. This allows for inherent parallel processing in computations. However, when concluding a calculation on a quantum computer, the resulting quantum-mechanical state needs to be measured to convert the calculation result into classical bits. Due to the superposition, the result of the measurement is not definite, but instead the values 0 and 1 of are obtained with a certain probability. The case of Eq. 6, |$\ket{0}$ is measured with a probability of ${|\alpha|}^{2}$ and |$\ket{1}$ with ${|\beta|}^{2}$. Gate-based quantum computers employ gates to perform operations on one or multiple qubits. Gates are unitary matrices. A sequence of multiple gates acting on one or more qubits is called a quantum circuit.
Despite those advantages, current quantum devices suffer from noise. We will now proceed to describe the quantum system and its interaction with the environment. Let us consider that the pure quantum system is described by a density matrix $\rho$. Thus, we can describe the incoherent noise, so the noise from the environment by a linear and completely trace preserving map $\varepsilon(\rho)$. Hereby a convenient strategy is to employ the Kraus operators, so a set of operators ${K_{i}}$, with the property: $\sum_{i}K_{i}^{\dagger}K_{i}\leq I$  [^1] Note that $A^{\dagger}$ describes the conjugate transpose of the matrix $A$.. With those operators, the quantum channel can be described as:
$\varepsilon(\rho)=\sum_{i}K_{i}^{\dagger}\rho K_{i}$ (7)
@@
  • $K_{0}=\sqrt{1-p}I$
  • $K_{1}=\sqrt{\frac{p}{3}}X$
  • $K_{2}=\sqrt{\frac{p}{3}}Y$
  • $K_{3}=\sqrt{\frac{p}{3}}Z$
@@
$I=\begin{pmatrix}1&0\\ 0&1\end{pmatrix},\;X=\begin{pmatrix}0&1\\ 1&0\end{pmatrix},\;Y=\begin{pmatrix}0&-i\\ i&0\end{pmatrix},\;Z=\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}$
@@ 2.2 Quantum Neural Network
In the setting of quantum computing, QNNs are realized trought what is known as a variational quantum algorithm (VQA) [3]. The execution of these algorithms is a blend of quantum and classical processes. On the one hand, the learning model is described entirely by a quantum circuit or integrates a quantum circuit at a certain point. The optimization on the other hand is managed using classical methods. A central element in this quantum algorithmic structure is the parametrized quantum circuit (PQC). Composed of unitary transformations, for example including rotation gates:
@@
these circuits can be adjusted via the parameters $\theta$ affecting the rotation gates. Regarding training, the loss function is computed using the true labels and the measurements of the quantum device. However, the task of parameter updates to minimize this loss is handled classically, primarily due to its efficient performance in this task.
@@
This strategy is quite attractive, since it provides an exponential reduction of qubits for encoding our datapoints. As a next step a circuit ansatz with the classically optimizable parameters is provided. For this ansatz, numerous options exist [12, 4, 34, 29]. We select the ansatz of Schuld et al. [29] to be architecture-wise in line with Du et al. [6]. The last part of the quantum neural network describes the measurement; the various classes are typically are encoded on different basis states. Here, a set of qubits is selected to be measured. The measured outcomes a assigned to class probabilities. With those probabilities, the loss of the network can be computed and the parameters can be optimized by a classical optimizer. Note that a computation of the gradient is possible on a quantum device, by the so called "parameter-shift" rule.
@@ 2.3 Quantum Adversarial Robustness
Definition 1 (Certified Robustness)
We say $f(x)\in\mathbb{R}^{n}$ is certifiable robust with perturbation budget $\epsilon$ if
[…]
@@
Typically the space of adversarial pertubations is defined by a Norm-Ball. This Norm-Ball on which PGD and FGSM attack the network is specified by selecting a specific process for $\Pi_{x+\mathcal{S}}(x^{\prime})$. When using the sign function, the examples will act on the $L_{\infty}$-Norm. The fact that QNNs are vulnerable to adversarial perturbations motivates the discussion about specific measures to counteract adversarial examples. In this work, we will utilize adversarial accuracy as in Def. 2 to describe the network’s vulnerability.
@@ 2.4 Adversarial Defense Strategies
Throughout recent years, multiple strategies have been developed to make models more robust against adversarial examples. In this context, we focused on two adversarial defense strategies: adversarial training and randomized smoothing
@@
Randomized smoothing, not only a method for robustness certification but also an adversarial training technique [27], involves adding noise to training data. This noise, typically Gaussian or Uniform with zero mean, mathematically expressed as $x=x+\delta$, with $\delta\sim\mathcal{N}(0,\sigma^{2}I)$ or $\delta\sim\mathcal{U}(-\sigma,\sigma)$, alters the training data. Unlike standard adversarial training that learns the worst-case decision boundary, randomized smoothing, guides the classifier to learn a smoothed decision boundary. This approach integrates Gaussian perturbations with adversarial examples, not just enhancing empirical robustness but also significantly boosting certifiable robust accuracy of the smoothed classifiers.
@@ 2.5 Quantum Certified Robustness
In the context of QML, the robustness bound is motivated by so called quantum differential privacy. Similarly towards classical differential privacy, quantum differential privacy assumes two datapoints $\sigma$ and $\rho$, separated by a small trace-distance: $\tau(\sigma,\rho)=\sqrt{1-\text{Tr}(\sigma,\rho)}$ and for all measurable sets $S$ it must hold that:
@@
Figure 2: Example Circuit with deploarization channel $\varepsilon_{p_{1}}$ and two final measurements
@@ 3 Robustness Analysis
As mentioned earlier, our goal is to analyze the robustness of QNNs and the effect of the bound provided by Du et al. [6]. Henceforth, we examined the following three research questions as relevant:
@@
To answer those questions, we utilize two datasets: Iris [7], which was motivated from the usage of Du et al. [6], and MNIST [16]. We train our models by the ansatz proposed by Schuld et al. [29] and encode our data on the quantum circuit using amplitude encoding [22].
@@ 3.1 Data-Preprocessing
For Iris, we followed the pre-processing of Du et al. [6], which resulted in the removal of the versicolor species and the petal_width feature. For the resulting dataset, we reported 100 feature vectors $x\in\mathbb{R}^{3}$, which were padded by a 00 and normalized by their respective $L_{2}$-Norm. Splitting  [^3] Splitting of the Iris dataset was done at random, utilizing sklearn’s train_test_split function. the dataset into a train and test set resulted in 60 training samples and 40 test samples. As a last step, we one-hot encoded class labels to make them compatible with the resulting measurement outcome from the quantum device.
@@
For the MNIST10 dataset, we rescaled the $28\times 28$ images into smaller images of the size $16\times 16$. Flattening the image resulted in a feature vector $x\in\mathbb{R}^{256}$, which can be realized with eight qubits. Next, we normalized the individual elements of the feature vector to be in the range $[0;1]$ and finally split the data into a training set of $48.000$ images, a validation set of $10.800$ images, and a test set of $1200$ images. To provide additional insight, specific classes of the dataset were subselected to mimic a binary classification task, MNIST2, and a four-class multiclassification task, MNIST4. For MNIST2, we used the classes for the digits one and nine. For MNIST4, we extended the present labels in MNIST2, providing overall the labels one, nine, seven, and three. MNIST2 consisted of $12.691$ datapoints, split  [^4] Splitting of MNIST2, MNIST4 and MNIST10 where done at random, utilizing sklearn’s train_test_split function twice once with with a test_size argument of 0.2 for training and test/validation set and once with a test_size of 0.1 for separating the test and validation set. into $10.152$ training images, $2285$ validation images, and $254$ testing images. Analougus MNIST4, consisting of $25.087$ images, was split into train-, test- and validation-set with the respective sizes $20.069$, $4516$, and $502$.

### 21 Dec 2023 v2

@@ 1 Introduction
The field of quantum computing has shown fast advancements in recent years, providing the possibility to speed up specific tasks. Hereby a variety of algorithms exists such as Shor’s algorithm to factor primes [35], Grover’s search to find an element in an unsorted list in $\mathcal{O}(\sqrt{n})$ [13] or the HHL-Algorithm to efficiently solve systems of linear equations [14]. In this context, quantum machine learning (QML) focuses onto providing the properties and speedups of quantum computing in the field of machine learning [39, 37, 33, 3]. First coined by Kak [17], the term quantum neural network (QNN) emerged as an ambitious effort to merge observations from the domain of neuroscience with the intriguing features inherent to quantum computations. Essentially, QNNs represent a specific type of Quantum Machine Learning (QML) models. Inspired by the layered approach of classical neural networks.
@@
As their classical counterpart, Lu et al. [23] highlighted that QNNs are not immune to adversarial attacks. Adversarial attacks involve carefully designed disturbances that can mislead machine learning models, causing them to produce incorrect predictions. Such attacks have long been a significant issue in traditional machine learning [4, 36, 12, 24]. The importance of investigating adversarial attacks on QNNs is reinforced by the no free lunch theorem [29]. This principle suggests that there is not a single model or algorithm universally optimal for all tasks, hinting at inherent vulnerabilities also in quantum-based models. Hence, finding these weak spots is crucial for the resilience and protection of quantum machine learning systems. A further obstacle is that current quantum hardware used to execute these models suffers from noise and decoherence times, which is the reason why Preskill [30] coined the term NISQ-Era for the current state in quantum computing. This means, that due the noise, limited connectivity and decoherence times, inherent in quantum devices, current algorithms can only be executed on smaller toy problems, providing a proof of concept, but suffering from not yet being applicable to practical problems.
Our Contribution
In this work, we explore the effect of noise, specifically depolarization noise, on the robustness of Quantum Neural Networks (QNNs), aiming to minimize compromises on accuracy and training data. We assumed the structure in Figure 1 as the most realistic scenario, so the input is perturbed on the classical side and then encoded on the quantum circuit. Furthermore, our research builds on the theoretical insights of Du et al. [9], which suggest that QNNs inherently possess an upper bound of robustness relative to the device’s depolarization noise. This upper bound arises because depolarizing noise, viewed as a stochastic Pauli-channel, effectively scrambles the state of the quantum device, thus blurring information. This blurring leads to a natural tradeoff between accuracy and robustness. We evaluate the effectiveness of this phenomenon by adding multiple depolarizing noise channels to a quantum neural network and comparing its performance with other adversarial training techniques commonly used in classical settings.
Our results indicate that depolarization noise does not improve adversarial robustness in multi-class classifiers in realistic settings. Additionally, we observed that increasing the number of classes diminishes both accuracy and overall robustness, with depolarization noise offering no significant enhancement.
2 Related Work
Since quantum neural networks are prone to adversarial attacks [22, 23], over recent years, multiple authors have tried to use quantum computing techniques to enhance adversarial robustness for both classical and quantum neural networks. As an example, Sahdev and Kumar [31] have utilized a quantum phase estimation to provide a quadratic speedup for randomized smoothing. However, because quantum phase estimation is not feasible on NISQ-Hardware and Sahdev and Kumar [31] requires a nested quantum phase estimation specifically for quantum neural networks, we did not use this method as a comparison in our studies. Gong et al. [11] proposed a vanishing gradient for introducing adversarial noise on a quantum circuit. Nevertheless, this method assumes that the data is already in a quantum state, subject to adversarial perturbation during the execution of the quantum circuit. Since we focused on encoding classical, adversarial perturbed data on the quantum circuit rather than assuming already encoded quantum states, we refrained from including Gong et al. [11] in our study. In the context of encoding classical data which is adversarially perturbed, Weber et al. [42] have provided a fundamental link between quantum hypothesis testing and binary classification. Their work also shows the connection between classical randomized smoothing and depolarization noise on a quantum circuit. Furthermore during the course of this study, Huang et al. [15] have provided a theoretical study on the fact that random noise in the form of rotational gates can improve the overall robustness. Additionally, Berberich et al. [1] have introduced a method for training weights to increase the Lipschitz bound of quantum neural networks, thereby enhancing their robustness. Considering specifically noise in quantum neural networks, Wang et al. [40] suggests that error mitigation of noise as well as noise aware training helps to improve the overall accuracy of the classifier.
@@ 3.1 Quantum information
$\psi=$ $\alpha\ket{0}+\beta\ket{1}\quad\text{with}$ $~{}{|\alpha|}^{2}+{|\beta|}^{2}=1.$ (1)
Unlike classical computers that use bits with values of either 0 or 1, quantum computers can simultaneously represent both states. This allows for inherent parallel processing in computations. However, when concluding a calculation on a quantum computer, the resulting quantum-mechanical state needs to be measured to convert the calculation result into classical bits. Due to the superposition, the result of the measurement is not definite, but instead the values 0 and 1 of are obtained with a certain probability. The case of Eq. 1, |$\ket{0}$ is measured with a probability of ${|\alpha|}^{2}$ and |$\ket{1}$ with ${|\beta|}^{2}$.
3.2 Depolarization Noise
Since quantum computers are currently in the noisy intermediate state era (NISQ era) [30], quantum devices typically suffer from noise. We can describe this noise (specifically incoherent noise) by a linear and completely trace preserving map. To make this more clear, let us consider that the pure quantum system is described by a density matrix $\rho$. Thus, we can describe the incoherent noise, so the noise from the environment by a linear and completely trace preserving map $\varepsilon(\rho)$. Hereby a convenient strategy is to employ the Kraus operators, so a set of operators ${K_{i}}$, with the property: $\sum_{i}K_{i}^{\dagger}K_{i}\leq I$  [^1] Note that $A^{\dagger}$ describes the conjugate transpose of the matrix $A$.. With those operators, the quantum channel can be described as:
$\varepsilon(\rho)=\sum_{i}K_{i}^{\dagger}\rho K_{i}.$ (2)
@@ 3.2 Depolarization Noise
$K_{0}=\sqrt{1-p}I$ $K_{1}=\sqrt{\frac{p}{3}}X$
$K_{2}=\sqrt{\frac{p}{3}}Y$ $K_{3}=\sqrt{\frac{p}{3}}Z$
@@
$I=\begin{pmatrix}1&0\\ 0&1\end{pmatrix}$ $X=\begin{pmatrix}0&1\\ 1&0\end{pmatrix}$
$Y=\begin{pmatrix}0&-i\\ i&0\end{pmatrix}$ $Z=\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}$
@@ 3.3 Quantum Neural Network
In the setting of quantum computing, QNNs are realized through what is known as a variational quantum algorithm (VQA) [6]. The execution of these algorithms is a blend of quantum and classical processes. On the one hand, the learning model is described entirely by a quantum circuit or integrates a quantum circuit at a certain point. The optimization on the other hand is managed using classical methods. A central element in this quantum algorithmic structure is the parametrized quantum circuit (PQC). Composed of unitary transformations, for example including rotation gates:
@@
$R_{x}(\theta)$ $=\begin{pmatrix}\cos(\frac{\theta}{2})&-i\sin(\frac{\theta}{2})\\ -i\sin(\frac{\theta}{2})&\cos(\frac{\theta}{2})\end{pmatrix}$ (11)
these circuits can be adjusted via the parameters $\theta$ affecting the rotation gates. Regarding training, the loss function is computed using the true labels and the measurements of the quantum device.
@@
This strategy is quite attractive, since it provides an exponential reduction of qubits for encoding our datapoints. As a next step a circuit ansatz with the classically optimizable parameters is provided. For this ansatz, numerous options exist [16, 8, 41, 34]. We select the ansatz of Schuld et al. [34] to be architecture-wise in line with Du et al. [9]. The last part of the quantum neural network describes the measurement; the various classes typically are encoded on different basis states. Here, a set of qubits is selected to be measured. The measured outcomes a assigned to class probabilities. With those probabilities, the loss of the network can be computed and the parameters can be optimized by a classical optimizer. Note that a computation of the gradient is possible on a quantum device, by the so called "parameter-shift" rule.
@@ 3.4 Quantum Adversarial Robustness
Quantum adversarial robustness examines how specially crafted attacks can weaken a quantum-based machine learning model, just as traditional neural networks can be fooled by adversarial inputs [12, 19, 25]. While there are other means of attacking a machine learning based system, we will focus on evasion attacks on the quantum machine learning model. In the context of classical machine learning, adversarial robustness refers to the ability of a machine learning to maintain a stable prediction when confronted with adversarial examples.
@@
Definition 1 (Adversarial Robustness)
Consider a classifier $f$ that maps an input $x$ to an output $y$. The adversarial robustness at input $x$ can be defined as the smallest amount of perturbation $\delta$ needed to change the model’s prediction:
$\operatorname{argmax}_{c\in\mathcal{Y}}f_{c}(x+\delta)\neq\operatorname{argmax}_{c\in\mathcal{Y}}f_{c}(x)\},$
where $\|\delta\|\leq\epsilon$ is bounded by a maximum $\epsilon$ budget and ${x}_{adv}=x+\delta$ denotes an adversarial example.
@@
Typically the space of adversarial perturbations is defined by a Norm-Ball. This Norm-Ball on which PGD and FGSM attack the network is specified by selecting a specific process for $\Pi_{x+\mathcal{S}}(x^{\prime})$. When using the sign function, the examples will act on the $L_{\infty}$-Norm. The fact that QNNs are vulnerable to adversarial perturbations motivates the discussion about specific measures to counteract adversarial examples. In this work, we will utilize adversarial accuracy as in Def. 2 to describe the network’s vulnerability.
@@ 3.5 Adversarial Defense Strategies
Over recent years, multiple strategies have been developed to make models more robust against adversarial examples. In this context, we focused on two adversarial defense strategies: adversarial training and randomized smoothing
@@
Randomized smoothing, not only a method for robustness certification but also an adversarial training technique [32], involves adding noise to training data. This noise, typically Gaussian or Uniform with zero mean, mathematically expressed as $x=x+\delta$, with $\delta\sim\mathcal{N}(0,\sigma^{2}I)$ or $\delta\sim\mathcal{U}(-\sigma,\sigma)$, alters the training data. Unlike standard adversarial training that learns the worst-case decision boundary, randomized smoothing guides the classifier to learn a smoothed decision boundary. This approach integrates Gaussian perturbations with adversarial examples, not just enhancing empirical robustness but also significantly boosting certifiable robust accuracy of the smoothed classifiers.
Definition 3 (Certified Robustness)
We say $f(x)\in\mathbb{R}^{n}$ is certifiable robust with perturbation budget $\epsilon>0$ if
[…]
Differently from adversarial robustness (Def. 1), certified robustness has a stronger assurance: it provides a formal guarantee that a model’s predictions will not change for any input within a specified range of perturbations.
@@ 3.6 Quantum Certified Robustness
In the context of QML, the robustness bound is motivated by so called quantum differential privacy. Similarly to classical differential privacy, quantum differential privacy assumes two datapoints $\sigma$ and $\rho$, separated by a small trace distance: $\tau(\sigma,\rho)=\sqrt{1-\text{Tr}(\sigma,\rho)}$ and for all measurable sets $S$ it must hold that:
@@
Figure 2: Example Circuit with depolarization channel $\varepsilon_{p_{1}}$ and two final measurements.
@@ 4 Robustness Analysis
As mentioned earlier, our goal is to analyze the robustness of QNNs and the effect of the bound provided by Du et al. [9]. Consequently, we examined the following three research questions, which we deemed relevant:
@@
To answer those questions, we utilized the dataset used by Du et al. [9] to outline the reproducability of his results. Secondly we used the MNIST [20], Breast Cancer Wisconsin (BC) [5] and Pima Indians Diabetes Database(PID) [5]. We wanted to estimated the performance of CIFAR10, by truancing it towards classes zero and nine. We train all models on the ansatz proposed by  Schuld et al. [34] and encode our data on the quantum circuit using amplitude encoding.
@@ 4.1 Data-Preprocessing
We will outline the data preprocessing done for all the datasets. Typically this process included dataset specific preprocessing techniques (normalization, feature engineering) combined with a split into a train, validation and test set or into a train and test set, depending on the size of the dataset. We performed this split the partition was varied among the datasets  [^3] Splitting of the datasets was done at random, utilizing sklearn’s train_test_split function. For multiple splits we separated the validation and test set from the validation set.
4.1.1 Iris
@@
For Iris, we followed the pre-processing of Du et al. [9], which resulted in the removal of the versicolor species and the petal_width feature. For the resulting dataset, we reported 100 feature vectors $x\in\mathbb{R}^{3}$, which were padded by a 00 and normalized by their respective $L_{2}$-Norm. Splitting the dataset into a train and test set resulted in 60 training samples and 40 test samples. As a last step, we one-hot encoded class labels to make them compatible with the resulting measurement outcome from the quantum device.
4.1.2 Pima Indians Diabetes (PID)
For the Pima Indians Diabetes dataset, we first dropped all nan or duplicate features from the dataset, next to overcome the present class imbalance, we removed overall $232$ datapoints from the majority class at random. We used the remaining dataset with $536$ datapoints and split them into a train set consisting of $375$ datapoints and a test set of $161$ datapoints. The train and test set was standard normalized ($x=\nicefrac{{x-x_{\text{mean}}}}{{x_{\text{std}}}}$) [^4] We used the StandardScaler from the sklearn package and afterwards rescaled to the interval $\left(0;1\right]$ to allow a comparison among other datasets.
4.1.3 Breast Cancer Wisconsin (BC)
For the Breast Cancer Wisconsin dataset, we dropped the nan features and duplicates as for the Pima Indians Dataset. This resulted in overall $683$ features. During hyperparameter training of the QNN, we figured that standard normalization hindered the convergence of the model, which was the reason why after splitting the dataset into a train set ($546$ datapoints) and test set ($137$ datapoints) we only rescaled the data into the interval $\left(0;1\right]$. It was not necessary to drop datapoints of the majority class, since the BC dataset was quite balanced from our point of view (we found a ratio of $65:35$, which we considered acceptable given the $\approx 90\%$ accuracy on the test set)
4.1.4 MNIST
@@
For the MNIST10 dataset, we rescaled the $28\times 28$ images into smaller images of the size $16\times 16$. Flattening the image resulted in a feature vector $x\in\mathbb{R}^{256}$, which can be realized with eight qubits. Next, we normalized the individual elements of the feature vector to be in the range $[0;1]$ and finally split the data into a training set of $48.000$ images, a validation set of $10.800$ images, and a test set of $1200$ images. To provide additional insight, specific classes of the dataset were subselected to mimic a binary classification task, MNIST2, and a four-class multiclassification task, MNIST4. For MNIST2, we used the classes for the digits one and nine. For MNIST4, we extended the present labels in MNIST2, providing overall the labels one, nine, seven, and three. MNIST2 consisted of $12.691$ datapoints, split into $10.152$ training images, $2285$ validation images, and $254$ testing images. Analougus MNIST4, consisting of $25.087$ images, was split into train-, test- and validation-set with the respective sizes $20.069$, $4516$, and $502$.
4.1.5 CIFAR2
For CIFAR10 [18], we droped all classes except zero and nine, upscaled the images from size $32\times 32$ to $64\times 64$, after this rescaling, we computed the histogram of orieanted gradients (HOG) using the scikit-image library [38]. For computing the HOG-features we used a $8\times 8$ cells with a inner block size of $1\times 1$. Overall this resulted in a feature vector $x\in\mathbb{R}^{576}$. We split the dataset into a train set of size $8000$ a validation set of size $1800$ and a test set of size $200$. We labeled this dataset CIFAR2, since we converted the multiclass classification task to a binary classification task.